# On Safety and Liveness Filtering Using Hamilton-Jacobi Reachability Analysis

### v1

@@ III-B Default Liveness (Safety) Controller from HJ Reachability
Along with the value function and the BRTs, reachability analysis also provides a liveness/safety controller for the system (referred to as the default liveness/safety reachability controller from here on). At state $x$ and time $t$, the default liveness reachability controller is given as:
$\pi_{\text{live}}^{*}(x,t)=\arg\min_{u\in\mathcal{U}}\max_{d\in\mathcal{D}}\nabla V(x,t)\cdot f(x,u,d).$ (9)
@@
$d^{*}(x,t)=\min_{u\in\mathcal{U}}\arg\max_{d\in\mathcal{D}}\nabla V(x,t)\cdot f(x,u,d).$ (10)
Intuitively, at any time step, $\pi_{\text{live}}^{*}$ (9) tries to maximally steer the system towards the target set. To understand this, recall that the value function for the livenss problem (4) $V(x,t)$ represents the minimum cost the system can achieve over the time horizon $[t,T]$, starting from state $x$ and time $t$ under worse case disturbance. In the liveness problem, we would like the system to reach a state of lower value. $\pi_{\text{live}}^{*}$ seeks to accomplish this objective by finding a control which steers the system in the direction of greatest decent on the value function (i.e. minimizing the dot product between $\dot{x}$ and the spatial gradient of the value function $\nabla V$). Moreover, it can be shown that as long as the system starts inside $\mathcal{G}_{\text{live}}(t)$, it is guaranteed to eventually reach the target set under $\pi_{\text{live}}^{*}$ despite the worst-case disturbance [4].
On the other hand, the default safety reachability controller’s goal is keeping the system as far away from the target set as possible, and the value function at state $x\in\mathcal{X}$ $V(x)$ is the minimum distance between the system and the target set over the trajectory starting at state $x$ under any control signals. Therefore, the optimal safety-preserving control at state $x$ should steers system towards a direction that is most aligned with spatial gradients of the value function:
$\pi_{\text{safe}}^{*}(x,t)=\arg\max_{u\in\mathcal{U}}\min_{d\in\mathcal{D}}\nabla V(x,t)\cdot f(x,u,d)$ (11)
As long as the system starts outside $\mathcal{G}_{\text{unsafe}}$, it is guaranteed to avoid the target set at all times under $\pi_{\text{safe}}^{*}$, despite the worst case disturbance.
Remark 1
[…]
@@
Running Example (Liveness). As discussed earlier, in the liveness problem, we are interested in computing $\mathcal{G}_{\text{live}}(t)$, along with associated value function. This requires solving the HJI-VI (6). Traditional methods compute a numerical solution of the HJI-VI over a state space grid [18]; however, these methods suffer from the curse of dimensionality and are not suitable for a 6D system. Instead, we use DeepReach [5] to compute the value function. Rather than solving the HJI-VI over a grid, DeepReach represents the value function as a sinusoidal neural network and learn a parameterized approximation of the value function. Thus, memory and complexity requirements for training scale with the value function complexity rather than the grid resolution. To train the neural network, DeepReach uses self-supervision on the HJI-VI itself. Ultimately, it takes as input a state $x$ and time $t$, and it outputs the value $V_{\beta}(x,t)$, where $\beta$ are the parameters of the NN. We refer interested readers to [5] for further details.
For this system, we use a sinusoidal neural network with three hidden layers, and 512 neurons per layer. The training took approximately two hours for 200 thousand epochs on an NVIDIA RTX 4090 GPU. Using scenario optimization [15], we verify the trained neural network by computing a high confidence error bound over the accuracy of the learned value function. A slice of the obtained BRT projected over the $YZ$-plane is shown in Fig. 1. As long as the system starts inside the teal region, it is guaranteed to reach the pink region (landing pad) under $\pi_{\text{live}}^{*}$. The system trajectory under $\pi_{\text{live}}^{*}$ for one such starting state is shown in dark blue in Fig. 1(a) and the corresponding control profiles are shown in Fig. 1(b), (c).
Figure 1: (a) BRT slice for a rocket landing problem in the $(y,z)$ states (all other states = 0) shown in teal. The target set $\mathcal{L}$ is shown in light pink rectangle. The trajectory followed by the rocket from its initial position (shown with the black rocket icon) under the default controller is shown in dark blue. (b) $u_{1}$ and (c) $u_{2}$ controller profiles generated by the default reachability controller displaying bang-bang behavior.
@@
However, this (a) results in a jittery, high-energy control profile, and (b) does not take into account any performance cost for the controller, which results in a very high total cost for this controller (black bar in Fig. 6).
Running Example (Safety). The safety problem is analogous to the liveness problem, except that the roles of disturbance and control are now reversed, as the control tries to steer the system away from the target set. For this example, we leverage the Level Set Toolbox [20] and its reachability helper library HelperOC [1] to compute a numerical approximation of the BRT. Level-Set methods solve the HJB-VI numerically over a uniformly discretized state space grid. We use a 4D grid of 81 x 81 x 41 x 21 points for our computations. The solution of the HJB-VI provides us with the value function, $V(x,\tau)$, which can be used to extract the $\mathcal{G}_{\text{unsafe}}$ similar to (8). In this case, the value function converges after 0.5 seconds; thus, we consider the time converged $\mathcal{G}_{\text{unsafe}}$. A 3D slice of $\mathcal{G}_{\text{unsafe}}$ for a fixed value of $\theta=0$ is presented in Fig.2. The expansion of the BRT can be seen in the reverse direction, as states oriented this way will inevitably collide with obstacles if they start too close. By definition of the BRT, as long as the system starts outside the blue region, it is guaranteed to remain safe under the default safety controller $\pi_{\text{safe}}^{*}$ (11).
Figure 2: A slice of the BRT for the safety example in the $(x,y,z)$ states for a fixed $\theta=0$. The system trajectory under the default safety controller is shown in dark blue and the corresponding controller profiles are shown in the inlets.
The system trajectory under $\pi_{\text{safe}}^{*}$ starting at a state outside the BRT and the corresponding control profiles are shown in Fig. 2. Once again, the system remains safe under the default safety controller as expected; however, the default safety controller steers the system as far away from the obstacles as possible, disregarding any other performance criterion the system might have (such as reaching the pink goal region in this case). Moreover, the control profile is jittery, which is typically undesirable for real-world applications.
@@ IV-A Characterizing the set of live controls
At state $x$ and time $t$, the set of all liveness-ensuring control inputs (abbreviated as live controls for the remainder of the paper) is given as:
$\mathcal{U}_{\text{live}}(x,t)$ =$=\{u:\exists\epsilon>0,$
$V\left(\xi_{x,t}^{u,d^{*}}(t+\delta),t+\delta\right)\leq 0~{}~{}\forall\delta\in[0,\epsilon]\},$ (12)
Here, $d^{*}$ is the optimal disturbance given by (10). Intuitively, $\mathcal{U}_{\text{live}}(x,t)$ (IV-A) is the set of all controls that instantaneously keep the next state of the system inside (or at the boundary of) $\mathcal{G}_{\text{live}}$, despite the worst case disturbance $d^{*}$. This ensures that the system can still reach the target set from the next state, thereby ensuring liveness recursively.
We will next use (IV-A) to characterize $\mathcal{U}_{\text{live}}(x,t)$. We assume that $x\notin\mathcal{L}$; otherwise, the system state is already within the target set and no further control action is needed.
@@
Assume that the value function $V(x,t)$ is differentiable for all $x$ and $t$. Let $x$ be any state such that $x\notin\mathcal{L}$. Then the set of live controls at $x$ is given as:
[…]
The proof of Lemma 1 is in the Appendix. Intuitively, (13) states that $\mathcal{U}_{\text{live}}(x,t)$ is the set of all admissible controls $\mathcal{U}$ when the system state is in the interior of $\mathcal{G}_{\text{live}}$. This makes sense as the system can apply any control and instantaneously remain inside $\mathcal{G}_{\text{live}}$ for a sufficiently small time step. On the other hand, when the system state is outside $\mathcal{G}_{\text{live}}$, liveness can’t be ensured despite the best control effort, resulting in an empty set. Finally, when the system state is at the boundary of $\mathcal{G}_{\text{live}}$, $\mathcal{U}_{\text{live}}(x,t)$ is given by the controls that ensure the total derivative of the value function, $\frac{d}{dt}V(x,t)$, to be 0 (i.e., the control inputs that keep the system state instantaneously at the boundary of $\mathcal{G}_{\text{live}}$).
@@
    For control and disturbance-affine system dynamics, the condition of $\frac{d}{dt}V(x,t)=0$ is linear in $u$. To see this, suppose the system dynamics are given by $f(x,u,d)=f_{1}(x)+f_{2}(x)u+f_{3}(x)d$. Its total derivative $\frac{d}{dt}V(x,t)$ is given by:
    $D_{t}V(x,t)+\nabla V(x,t)\cdot f(x,u,d^{*})$
    $=D_{t}V(x,t)+\nabla V(x,t)\cdot(f_{1}(x)+f_{2}(x)u+f_{3}(x)d^{*})$
@@
    We can observe that $\frac{d}{dt}V(x,t)$ is now a linear equation of $u$, where $\alpha=D_{t}V(x,t)+\nabla V(x,t)\cdot(f_{1}(x)+f_{3}(x)d^{*})$ and $\beta=\nabla V(x,t)\cdot f_{2}(x)$. Thus, $\mathcal{U}_{\text{live}}(x,t)$ is given by a hyperplane in $\mathcal{U}$ (a point if the control space is one-dimensional), whenever the system state is at the boundary of $\mathcal{G}_{\text{live}}(t)$. As we will see later, this will make the liveness filter design particularly computationally efficient for control-affine systems.
@@
    Control inputs given by the default liveness controller (9) is always contained within $\mathcal{U}_{\text{live}}(x,t)$. The following corollary formalizes this result.
@@
[…]
@@
The real utility of (13) is that it can potentially provide more than one liveness-ensuring control inputs, whereas the default liveness controller can only provide one at a given state $x$ and time $t$. We now formally establish that applying any control inputs from $\mathcal{U}_{\text{live}}(x,t)$ is sufficient to ensure that the system eventually reaches the target set within in the time horizon $[t,T]$.
@@
[…]
Lemma 2 effectively allows us to use any subset of $\mathcal{U}_{\text{live}}(x,t)$ for the projection of the nominal controller, while still maintaining liveness. This suggests the following general structure for liveness filters:
@@
$\begin{split}\pi(x,t)=\arg\min_{u}&\quad h(u,\pi_{\text{nom}}(x,t))\\ s.t.~{}&u\in\widetilde{\mathcal{U}}(x,t),\text{~{}with~{}}\widetilde{\mathcal{U}}(x,t)\subseteq\mathcal{U}_{\text{live}}(x,t)\end{split}$ (14)
In (14), we refer to $h$ as the projection operator and $\widetilde{\mathcal{U}}$ as the projection set. Since $\widetilde{\mathcal{U}}\subseteq\mathcal{U}_{\text{live}}$, the above filter makes sure that $\pi(x,t)\in\mathcal{U}_{\text{live}}(x,t)$, thereby ensuring system liveness at all times (by Lemma 2). Nevertheless, different choices of $h$ and $\hat{\mathcal{U}}$ will lead to different trade-offs between performance and liveness.
To make sure that the optimization problem (14) can be solved in a computationally efficient manner, it is often desirable to use a projection operator and set that are convex in $u$ for a given $x$ and $t$. A particularly popular choice in the literature for $h$ is the $l_{2}$ distance from the nominal controller, i.e., $h:=||u-\pi_{\text{nom}}(x,t)||_{2}^{2}$. For the remainder of this section, we use this $l_{2}$ projection operator and choose three different $\widetilde{\mathcal{U}}$ that result in particularly interesting liveness filters. We will conclude the section by comparing the resultant filters.
@@ IV-B Least Restrictive Filter
$\pi(x,t)=\begin{cases}\pi_{\text{nom}}(x,t)&\quad\text{ if }V(x,t)<0\\ \pi_{\text{live}}^{*}(x,t)&\quad\text{ if }V(x)=0\end{cases}$ (15)
The least restrictive (LR) filter follows the nominal controller when the system state $x$ is in the interior of $\mathcal{G}_{\text{live}}(t)$ (i.e. $V(x,t)<0$), and it takes corrective actions given by $\pi^{*}_{\text{live}}(x,t)$ when the system is on the boundary of or at the risk of exiting $\mathcal{G}_{\text{live}}(t)$. This ensures that the system never exits $\mathcal{G}_{\text{live}}(t)$. The controller in (15) is least restrictive in the sense that it follows the given nominal controller that can optimize criteria besides liveness, and only interferes when the system is at the risk of breaching liveness. Consequently, the system can optimize other objectives while progressing towards the goal, but such freedom is not afforded by $\pi_{\text{live}}^{*}(x,t)$ as its sole objective is to ensure system’s liveness.
@@
$\widetilde{\mathcal{U}}(x,t)=\begin{cases}\mathcal{U}&\quad\text{ if }V(x,t)<0\\ \{\pi_{\text{live}}^{*}(x,t)\}&\quad\text{ if }V(x)=0\end{cases}$ (16)
Thus, $\widetilde{\mathcal{U}}$ is the set of all permissible controls whenever the system state is inside $\mathcal{G}_{\text{live}}(t)$ and is given by a singleton when it is at the BRT boundary. It is easy to verify that $\widetilde{\mathcal{U}}(x,t)\subset\mathcal{U}_{\text{live}}(x,t)$ for all $x$ and $t$. Thus, the system will always remain live under the LR filter.
Running Example (Liveness). We now demonstrate the LR filter (15) on the running example. The corresponding system trajectory and the control profile (for $u_{1}$) are shown in purple in Fig. 3 and 4, respectively. As expected, the LR filter chooses between $\pi^{*}_{\text{live}}$ and $\pi_{\text{nom}}$ to steer the system towards the target set. However, as discussed earlier, using $\pi^{*}_{\text{live}}$ results in bang-bang behavior (as evident from the spikes in the control profile) and high control energy (purple bar in Fig. 6). In the next subsection, we propose a smoother version of the LR filter to overcome this challenge.
Figure 3: System trajectories under the nominal and filtered liveness controllers for the rocket landing system.

### v2

@@ III-B Default Liveness (Safety) Controller from HJ Reachability
Along with the value function and the BRTs, reachability analysis also provides a liveness/safety controller for the system (referred to as the default liveness/safety reachability controller from here on). At state $\mathbf{x}$ and time $t$, the default liveness reachability controller is given as:
$\pi_{\text{live}}^{*}(\mathbf{x},t)=\arg\min_{u\in\mathcal{U}}\max_{d\in\mathcal{D}}\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d)$ (9)
@@
$d^{*}(\mathbf{x},t)=\min_{u\in\mathcal{U}}\arg\max_{d\in\mathcal{D}}\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d)$ (10)
Intuitively, at any time step, $\pi_{\text{live}}^{*}(\mathbf{x},t)$ tries to maximally steer the system state towards the target set (we omit the state and time dependencies of the optimal controller and disturbance onwards for compactness). To understand this, recall that the value function for the liveness problem (4), $V(\mathbf{x},t)$ represents the minimum cost the system can achieve over the time horizon $[t,T]$, starting from state $\mathbf{x}$ and time $t$ under worst case disturbance. In the liveness problem, we would like the system to reach a state of lower value. $\pi_{\text{live}}^{*}$ seeks to accomplish this objective by finding a control which steers the system in the direction of greatest decent on the value function (i.e. minimizing the dot product between $\dot{\mathbf{x}}$ and the spatial gradient of the value function $\nabla V$). Moreover, it can be shown that as long as the system starts inside $\mathcal{G}_{\text{live}}(t)$, it is guaranteed to eventually reach the target set under $\pi_{\text{live}}^{*}$ despite the worst-case disturbance [3]. This is analogous to the safety case, with the difference that the controller would try to steer the system away from the target set.
@@
Running Example (Liveness). As discussed earlier, in the liveness problem, we are interested in computing $\mathcal{G}_{\text{live}}(t)$, along with associated value function. This requires solving the HJI-VI (6). Traditional methods compute a numerical solution of the HJI-VI over a state space grid [19]; however, these methods suffer from the curse of dimensionality and are not suitable for a 6D system. Instead, we use DeepReach [4] to compute the value function. Rather than solving the HJI-VI over a grid, DeepReach represents the value function as a sinusoidal neural network and learn a parameterized approximation of the value function. Thus, memory and complexity requirements for training scale with the value function complexity rather than the grid resolution. To train the neural network, DeepReach uses self-supervision on the HJI-VI itself. Ultimately, it takes as input a state $\mathbf{x}$ and time $t$, and it outputs the value $V_{\beta}(\mathbf{x},t)$, where $\beta$ are the parameters of the NN. We refer interested readers to [4] for further details.
For this system, we use a sinusoidal neural network with three hidden layers, and 512 neurons per layer. The training took approximately two hours for 20000 epochs on an NVIDIA RTX 4090 GPU. Using scenario optimization [16], we verify the trained neural network by computing a high confidence error bound over the accuracy of the learned value function. A slice of the obtained BRT projected over the $YZ$-plane is shown in Fig. 1. As long as the system starts inside the teal region, it is guaranteed to reach the pink region (landing pad) under $\pi_{\text{live}}^{*}$. The system trajectory under $\pi_{\text{live}}^{*}$ for one such starting state is shown in dark blue in Fig. 1(a) and the corresponding control profiles are shown in Fig. 1(b), (c).
Figure 1: (a) BRT slice for a rocket landing problem in the $(y,z)$ states (all other states = 0) shown in teal. The target set $\mathcal{L}$ is shown in light pink rectangle. The trajectory followed by the rocket from its initial position (shown with the black rocket icon) under the default controller is shown in dark blue. (b) $u_{y}$ and (c) $u_{z}$ controller profiles generated by the default reachability controller displaying bang-bang behavior.
@@
However, as we will show in later sections, this approach results in jittery, high-energy control profiles that do not take into account any performance criteria, which results in a very high total cost for this control.
@@ IV-A Characterizing the set of live controls
At state $\mathbf{x}$ and time $t$, the set of all liveness-ensuring control inputs (abbreviated as live controls for the remainder of the paper) is given as:
$\mathcal{U}_{\text{live}}(\mathbf{x},t)$ =$=\{u:\exists\epsilon>0,$
$V\left(\xi_{\mathbf{x},t}^{u,d^{*}}(t+\delta),t+\delta\right)\leq 0~{}~{}\forall\delta\in[0,\epsilon]\}$ (11)
Here, $V$ represents the value function obtained using the HJI-VI in (6). $d^{*}$ is the optimal disturbance given by (10). Intuitively, $\mathcal{U}_{\text{live}}(\mathbf{x},t)$ (IV-A) is the set of all controls that instantaneously keep the next state of the system inside (or at the boundary of) $\mathcal{G}_{\text{live}}$, despite the worst case disturbance $d^{*}$. This ensures that the system can still reach the target set from the next state, thereby ensuring liveness recursively.
We will next use (IV-A) to characterize $\mathcal{U}_{\text{live}}(\mathbf{x},t)$. We assume that $\mathbf{x}\notin\mathcal{L}$; otherwise, the system state is already within the target set and no further control action is needed.
@@
Assume that the value function $V(\mathbf{x},t)$ is differentiable for all $\mathbf{x}$ and $t$. Let $\mathbf{x}$ be any state such that $\mathbf{x}\notin\mathcal{L}$ and $d^{*}\in\mathcal{D}$ the optimal disturbance that maximally decreases liveness. Then the set of live controls at $\mathbf{x}$ is given as:
$\mathcal{U}_{\text{live}}(\mathbf{x},t)=\left\{\begin{array}[]{l}\mathcal{U}\quad\text{ if }V(\mathbf{x},t)<0\\ \\ \{u\in\mathcal{U}:\\ D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d^{*})=0\}\\ \quad\quad\quad\quad\quad\quad\text{ if }V(\mathbf{x},t)=0\\ \\ \emptyset\quad\text{ if }V(\mathbf{x},t)>0\end{array}\right.$ (12)
The proof of Lemma 1 is in the Appendix. Intuitively, (12) states that $\mathcal{U}_{\text{live}}(\mathbf{x},t)$ is the set of all admissible controls $\mathcal{U}$ when the system state is in the interior of $\mathcal{G}_{\text{live}}$. This makes sense as the system can apply any control and instantaneously remain inside $\mathcal{G}_{\text{live}}$ for a sufficiently small time step. On the other hand, when the system state is outside $\mathcal{G}_{\text{live}}$, liveness can’t be ensured despite the best control effort, resulting in an empty set. Finally, when the system state is at the boundary of $\mathcal{G}_{\text{live}}$, $\mathcal{U}_{\text{live}}(\mathbf{x},t)$ is given by the controls that ensure the total derivative of the value function, $\frac{d}{dt}V(\mathbf{x},t)$, to be 0 (i.e., the control inputs that keep the system state instantaneously at the boundary of $\mathcal{G}_{\text{live}}$).
@@
    For control and disturbance-affine system dynamics, the condition of $\frac{d}{dt}V(\mathbf{x},t)=0$ is linear in $u$. To see this, suppose the system dynamics are given by $f(\mathbf{x},u,d)=f_{1}(\mathbf{x})+f_{2}(\mathbf{x})u+f_{3}(\mathbf{x})d$. Its total derivative $\frac{d}{dt}V(\mathbf{x},t)$, under the optimal disturbance $d^{*}$, is given by:
    $D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot f(\mathbf{x},u,d^{*})$
    $=D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot(f_{1}(\mathbf{x})+f_{2}(\mathbf{x})u+f_{3}(\mathbf{x})d^{*})$
@@
    We can observe that $\frac{d}{dt}V(\mathbf{x},t)$ is now a linear equation of $u$, where $\alpha=D_{t}V(\mathbf{x},t)+\nabla V(\mathbf{x},t)\cdot(f_{1}(\mathbf{x})+f_{3}(\mathbf{x})d^{*})$ and $\beta=\nabla V(\mathbf{x},t)\cdot f_{2}(\mathbf{x})$. Thus, $\mathcal{U}_{\text{live}}(\mathbf{x},t)$ is given by a hyperplane in $\mathcal{U}$ (a point if the control space is one-dimensional), whenever the system state is at the boundary of $\mathcal{G}_{\text{live}}(t)$. As we will see later, this will make the liveness filter design computationally efficient for control-affine systems.
@@
    Control inputs given by the default liveness controller (9) is always contained within $\mathcal{U}_{\text{live}}(\mathbf{x},t)$. The following corollary formalizes this result.
@@
    $\pi_{\text{live}}^{*}(\mathbf{x},t)\in\mathcal{U}_{\text{live}}(\mathbf{x},t)\ \forall\mathbf{x}\in\mathcal{G}_{\text{live}}$.
@@
The real utility of (12) is that it can potentially provide more than one liveness-ensuring control inputs, whereas the default liveness controller can only provide one at a given state $\mathbf{x}$ and time $t$. We now formally establish that applying any control inputs from $\mathcal{U}_{\text{live}}(\mathbf{x},t)$ is sufficient to ensure that the system will eventually reach the target set within the established time horizon $[t,T]$.
@@
Suppose the system starts inside $\mathcal{G}_{\text{live}}$ at time $t$. If the system applies control $u(\tau)\in\mathcal{U}_{\text{live}}(\mathbf{x}(\tau),\tau)\ \forall\tau\in[t,T]$, then $\exists s\in[t,T]$ such that $\mathbf{x}(s)\in\mathcal{L}$.
Lemma 2 effectively allows us to use any subset of $\mathcal{U}_{\text{live}}(\mathbf{x},t)$ for the projection of the nominal controller, while still maintaining liveness. This suggests the following general structure for liveness filters:
@@
$\begin{split}\pi(\mathbf{x},t)=\arg\min_{u}&\quad h(u,\pi_{\text{nom}}(\mathbf{x},t))\\ s.t.~{}&u\in\widetilde{\mathcal{U}}(\mathbf{x},t),\text{~{}with~{}}\widetilde{\mathcal{U}}(\mathbf{x},t)\subseteq\mathcal{U}_{\text{live}}(\mathbf{x},t)\end{split}$ (13)
In (13), we refer to $h$ as the projection operator and $\widetilde{\mathcal{U}}$ as the projection set. Since $\widetilde{\mathcal{U}}\subseteq\mathcal{U}_{\text{live}}$, the above filter makes sure that $\pi(\mathbf{x},t)\in\mathcal{U}_{\text{live}}(\mathbf{x},t)$, thereby ensuring system liveness at all times (by Lemma 2). Nevertheless, different choices of $h$ and $\widetilde{\mathcal{U}}$ will lead to different trade-offs between performance and liveness.
To make sure that the optimization problem (13) can be solved in a computationally efficient manner, it is often desirable to use a projection operator and set that are convex in $u$ for a given $\mathbf{x}$ and $t$. A particularly popular choice in the literature for $h$ is the $l_{2}$ distance from the nominal controller, i.e., $h:=||u-\pi_{\text{nom}}(\mathbf{x},t)||_{2}^{2}$. For the remainder of this section, we use this $l_{2}$ projection operator and choose three different $\widetilde{\mathcal{U}}$ that result in particularly interesting liveness filters. We will conclude the section by comparing the resultant filters.
@@ IV-B Least Restrictive Filter
$\pi(\mathbf{x},t)=\begin{cases}\pi_{\text{nom}}(\mathbf{x},t)&\quad\text{ if }V(\mathbf{x},t)<0\\ \pi_{\text{live}}^{*}(\mathbf{x},t)&\quad\text{ if }V(\mathbf{x})=0\end{cases}$ (14)
The least restrictive (LR) filter follows the nominal controller when the system state $\mathbf{x}$ is in the interior of $\mathcal{G}_{\text{live}}(t)$ (i.e. $V(\mathbf{x},t)<0$), and it takes corrective actions given by $\pi^{*}_{\text{live}}(\mathbf{x},t)$ when the system is on the boundary of or at the risk of exiting $\mathcal{G}_{\text{live}}(t)$. This ensures that the system never exits $\mathcal{G}_{\text{live}}(t)$. The controller in (14) is least restrictive in the sense that it follows the given nominal controller that can optimize criteria besides liveness, and only interferes when the system is at the risk of breaching liveness. Consequently, the system can optimize other objectives while progressing towards the goal, but such freedom is not afforded by $\pi_{\text{live}}^{*}(\mathbf{x},t)$ as its sole objective is to ensure system’s liveness.
@@
$\widetilde{\mathcal{U}}(\mathbf{x},t)=\begin{cases}\mathcal{U}&\quad\text{ if }V(\mathbf{x},t)<0\\ \{\pi_{\text{live}}^{*}(\mathbf{x},t)\}&\quad\text{ if }V(\mathbf{x},t)=0\end{cases}$ (15)
Thus, $\widetilde{\mathcal{U}}$ is the set of all permissible controls whenever the system state is inside $\mathcal{G}_{\text{live}}(t)$ and is given by a singleton when it is at the BRT boundary. It is easy to verify that $\widetilde{\mathcal{U}}(\mathbf{x},t)\subset\mathcal{U}_{\text{live}}(\mathbf{x},t)$ for all $\mathbf{x}$ and $t$. Thus, the system will always remain live under the LR filter.
Figure 2: System trajectories under the nominal and filtered liveness controllers for the rocket landing system. Pink region represents the target set.
Running Example (Liveness). We now demonstrate the LR filter (14) on the running example. The corresponding system trajectory and the control profile (for $u_{1}$) are shown in purple in Fig. 2 and Fig. 3, respectively. As expected, the LR filter chooses between $\pi^{*}_{\text{live}}$ and $\pi_{\text{nom}}$ to steer the system towards the target set. However, as discussed earlier, using $\pi^{*}_{\text{live}}$ results in bang-bang behavior (as evident from the spikes in the control profile for the LR filter in Fig. 3) and high control energy. In the next subsection, we propose a smoother version of the LR filter to overcome this challenge.